# Gabor systems with Hermite functions of order $n$ and oversampling greater than $n+1$ which are not frames

### v1

@@ Abstract.
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/P33217].
@@ 1. Introduction
Recently, Gabor frames with Hermite functions and oversampling rate $n+1$ were found by Faulhuber, Shafkulovska, and Zlotnikov [2]. On the other side, simple conditions for the frame property of Gabor systems with Hermite functions are to be ruled out eventually, due to the results of Lemvig [10] or Horst, Lemvig, and Videbaek [8].
In this work, we are concerned with the following question: can the sufficiency condition of Gröchenig and Lyubarskii be extended beyond the lattice setting? The answer is NO!
@@ 2. Notation
In most parts of this article we will follow the notation from the book of Gröchenig [5]. The Hilbert space of square-integrable functions on the real line is denoted by $L^{2}({\mathbb{R}})$. For technical reasons, it will sometime be necessary to restrict ourselves to a suitable dense subspace, in which case we will also write that a function is suitable. The statement is then valid for the Hilbert space by the density of the suitable function space, with the constraint that some results will only hold almost everywhere and not pointwise.
The modulation space $M^{1}({\mathbb{R}})$, also known as the Feichtinger algebra $S_{0}({\mathbb{R}})$ [3], is a suitable function space. In some situations, functions from the (slightly larger) Wiener space of continuous functions $W_{0}({\mathbb{R}})$ with Fourier transform also contained in $W_{0}({\mathbb{R}})$ suffice. In any case, our results will be for Hermite functions, which are contained in the Schwartz space $\mathcal{S}({\mathbb{R}})$ and the reader may consider their preferred function space as suitable.
@@
Moreover, we will use the unitary dilation operator $\mathcal{D}_{a}$, $a>0$, which acts on suitable functions by the rule
@@
This clarifies the basic essentials. In § 2.1 we clarify our normalization for Hermite functions, which is the same as in the book of Folland [4]. Time-frequency shifts and multi-window Gabor systems are introduced in § 2.2, following the book of Gröchenig [5]. Also, the notion of a periodic structure (the union of finitely many shifted copies of one lattice) is discussed there. We then introduce the (parameter free) Zak transform in § 2.3, which maps functions from ${\mathbb{R}}$ to quasi-periodic functions with period lattice ${\mathbb{Z}}^{2}$. This is again covered by the book of Gröchenig [5]. The reader familiar with the concepts and notation may directly jump to the results in § 3.
@@ 2.1. Hermite functions
We consider Hermite functions as defined in [4]. The $n$-th order Hermite function is the $n$-th Hermite polynomial $H_{n}$ (subject to our normalization) multiplied with a Gaussian function $\phi(t)=e^{-\pi t^{2}}$. More precisely, we have
@@ A Gabor system in $L^{2}({\mathbb{R}})$ is a function system of the type
(2.8) $\pi(z)=M_{\omega}T_{x},\ z=(x,\omega)\quad T_{x}f(t)=f(t-x),\quad M_{\omega}f(t)=e^{2\pi i\omega t}.$
@@
We refer to [5, Chap. 9.1] for more details. The important part for us here is the fact that the composition of two time-frequency shifts is (up to a phase factor) again a time-frequency shift by the sum of the co-ordinates.
@@
We remark that we always assume $\Gamma$ to be relatively separated, i.e.,
(2.13) $\min_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\Gamma\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}(|\gamma_{1}-\gamma_{2}|)\geq\delta>0.$
In this case, for $g\in M^{1}({\mathbb{R}})$ the upper frame bound is always finite and we only need to care about the strict positivity of the lower frame bound.
@@
(2.14) $\mathfrak{G}_{M}(g_{m},\Gamma)=\bigcup_{m=1}^{M}\mathcal{G}(g_{m},\Gamma)$
@@
(2.15) $S_{\mathfrak{G}_{M}(g_{m},\Gamma)}=\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},\Gamma)}=\sum_{m=1}^{M}\sum_{\gamma\in\Gamma}\langle f,\pi(\gamma)g_{m}\rangle\,\pi(\gamma)g_{m}.$
The windows $g_{m}$ may be completely different, but we are interested in the case where they are simply shifted copies $\pi(z_{m})g$ of a single function $g$. A direct computation shows that the multi-window frame operator is also the associated frame operator of a Gabor system over a periodic structure $\Gamma_{M}(\Lambda,\{z_{m}\}_{m=1}^{M})$ with window $g$. The periodic structure is
(2.16) $\Gamma_{M}(\Lambda,\{z_{m}\}_{m=1}^{M})=\bigcup\limits_{m=1}^{M}(\Lambda+z_{m})$
To simplify notation, we will only write $\Gamma_{M}(\Lambda)$, keeping in mind that we still need to choose the set $\{z_{1},\ldots,z_{m}\}$. While $\Gamma_{M}(\Lambda)$ is still $\Lambda$-periodic it is not necessarily a lattice. The density, or (over-)sampling rate, of a periodic structure is given by
(2.17) $\rho(\Gamma_{M}(\Lambda))=\frac{M}{{\rm vol}({\mathbb{R}}^{2}/\Lambda)}.$
Assume $g$ is suitably nice (so the following series converges unconditionally), then
(2.18) $S_{\mathfrak{G}_{M}(g_{m},\Gamma_{M}(\Lambda))}f$ $=\sum_{m=1}^{M}\sum_{\lambda\in\Lambda}\langle f,\pi(\lambda)\pi(z_{m})g\rangle\,\pi(\lambda)\pi(z_{m})g$
@@
The phase factor, which appears due to (2.9), also appears as complex conjugate in the above calculation and, hence, cancels. The last detail we collect is the unitary equivalence of the Gabor systems $\mathcal{G}(g,\Gamma)$ and $\mathcal{G}(\mathcal{D}_{a}g,D_{a}\Gamma)$, where
@@
First, for $z\in{\mathbb{R}}^{2}$, we note that $\mathcal{D}_{a}\,\pi(z)\,\mathcal{D}_{a}^{-1}=\pi(D_{a}z)$, which can be checked by a simple calculation with a suitable function. From this we get
@@
In the above computations we also used the unconditional convergence of the series (because $g$ is suitable, so we have a Bessel system, i.e., a finite upper frame bound) and the fact that $\mathcal{D}_{a}$ is unitary. In particular, the two Gabor systems have the same optimal frame bounds.
@@ 2.3. The Zak transform
Thus, $\mathcal{Z}f(x,\omega)$ is quasi-periodic with respect to the integer lattice ${\mathbb{Z}}^{2}$ and, thus, completely determined by the values in $[0,1)^{2}$. Next, we note that a time-frequency shift of the function basically results in a shift of the Zak transform. Due to (2.9), we obtain
@@
the phase factor in (2.27) appears then due to the commutation relation (2.9). We also note that the Zak transform extends to a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}([0,1]^{2})$. For the details of this fact we refer to [5, Thm. 8.2.3].
We close the survey on the Zak transform with listing zeros which occur due to the parity of $f$. For more details we refer to [9, § 5] or [8, § 4]. We have (see Fig. 1), for suitable $f$,
  • if $f(t)=f(-t)$, so $f$ is an even function, then
@@
The condition that $f$ is suitable cannot be dropped here. It implies that $\mathcal{Z}f$ is continuous. A remarkable property of the Zak transform is that, if $\mathcal{Z}f$ is continuous, then it has at least a zero in $[0,1)^{2}$. The indicator function $\chi_{[0,1]}$ is not suitable. It is not continuous, hence, in neither of the mentioned suitable function spaces and $|\mathcal{Z}\chi_{[0,1]}|=1$ almost everywhere.
@@ 3. Results
Let $\{z_{1},\ldots,z_{M}\}\subset[0,1)^{2}$ be distinct zeros of the Zak transform of $g$, i.e., $\mathcal{Z}g(z_{m})=0$, $m=1,\ldots,M$. Then, the multi-window Gabor system $\mathfrak{G}_{M}(\pi(z_{m})g,{\mathbb{Z}}^{2})$ is not a frame. Equivalently, the Gabor system $\mathcal{G}\left(g,\bigcup_{m=1}^{M}({\mathbb{Z}}^{2}+z_{m})\right)$ is not a frame.
@@
To keep this article self-contained, we repeat a standard proof, as given in the book of Gröchenig [5, Chap. 8.3], that the Zak transform diagonalizes the frame operator. More precisely, for a mulit-window Gabor system we will show (cf. [5, eq. (8.18)]) that
@@
(3.2) $\mathcal{Z}\left(\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right)=\sum_{m=1}^{M}\left(\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right),$
@@
(3.5) $=\left(\sum_{(k,l)\in{\mathbb{Z}}^{2}}\langle\mathcal{Z}f,\mathcal{Z}(M_{l}T_{k}g_{m})\rangle_{L^{2}([0,1]^{2})}\,e^{2\pi i(k\omega-lx)}\right)\,\mathcal{Z}g_{m}(x,\omega)$
We have a look at the inner product. Using (3.3) we obtain
@@
These are Fourier coefficients of $\mathcal{Z}f\overline{\mathcal{Z}g_{m}}$ and the whole expression in the parentheses in (3.5) is simply a Fourier series of this expression. Thus,
(3.7) $\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\left(\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}\right)\mathcal{Z}f\quad\Longleftrightarrow\quad\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}\mathcal{Z}^{-1}=\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}$
@@
(3.9) $\sum_{m=1}^{M}|\mathcal{Z}(\pi(z_{m})g)(z)|^{2}=\sum_{m=1}^{M}|\mathcal{Z}(g)(z+z_{m})|^{2}\quad z=(x,\omega).$
@@
forcing the lower frame bound to vanish and the result is proved. ∎
At this point we remark, that there are several versions of the Zak transform. For example, in [8] or [10] the following version is used (re-call $\mathcal{D}_{a}$ from (2.3))
@@
This can be used to study Gabor systems of the form $\mathcal{G}(\mathcal{D}_{a}^{-1}g,{\mathbb{Z}}^{2})$. By unitary equivalence, this is the same as studying properties of $\mathcal{G}(g,D_{a}{\mathbb{Z}}^{2})$, so a Gabor system with window g over the rectangular lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. On the other hand, there is the following version, which is for example presented in [9]. Setting $z=(x,\omega)\in{\mathbb{R}}^{2}$, we can write this version as
(3.12) $\mathcal{Z}_{a}f(z)$ $=\sqrt{a}\sum_{k\in{\mathbb{Z}}}f(ak-x)e^{2\pi iak\omega}$
(3.13) $=\sum_{k\in{\mathbb{Z}}}\mathcal{D}_{a}^{-1}\pi(z)f(k)=\sum_{k\in{\mathbb{Z}}}\pi(D_{a}^{-1}z)\mathcal{D}_{a}^{-1}f(k)$
@@
The essential properties from $\mathcal{Z}$ carry over to $\mathcal{Z}_{a}$, but the main difference is that $\mathcal{Z}_{a}$ is quasi-periodic with respect to the lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. Hence, we have that $\mathcal{Z}_{a}$ is determined by its values on $[0,a)\times[0,\frac{1}{a})$ and it is a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}(D_{a}[0,1]^{2})$. This version of the Zak transform, combined with the techniques presented in the proof above, has long been used to study Gabor systems over rectangular lattices of the form $a{\mathbb{Z}}\times b{\mathbb{Z}}$ with $(ab)^{-1}\in{\mathbb{N}}$. The technique has also been expanded to the case that $(ab)^{-1}=\frac{p}{q}\in{\mathbb{Q}}$, leading to determining whether certain $q\times q$ matrices have full rank (see [5, Chap. 8.3]). Having a rectangular lattice as an index set is quite restrictive, as the shifts $\pi(\xi_{n})$ must possess very restricted symmetries. This is discussed to some extent in [8, § 3]. The freedom of allowing arbitrary periodic configurations leads to the following result.
Corollary 3.2.
@@
The fact that $g\in M^{1}({\mathbb{R}})$ guarantees that we have a finite upper frame bound. As the zero set of $\mathcal{Z}g$ is discrete and finite by assumption, the probability that $K$ i.i.d. randomly picked picking points from $[0,1)^{2}$ all belong to the zero set of $\mathcal{Z}g$ is zero. ∎
@@
(3.16) $\Gamma_{n+1}({\mathbb{Z}}^{2},\{z_{\ell}\}_{\ell=1}^{n+1})=\bigcup_{\ell=1}^{n+1}({\mathbb{Z}}^{2}+z_{\ell})\text{ is a lattice.}$.

### v2

@@ Abstract.
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/P33217]. The author thanks Karlheinz Gröchenig for feedback on the initial draft of the paper. The author gratefully acknowledges the comments from the anonymous referees.
@@ 1. Introduction
In this work, we answer the following question, which has been around for about 20 years, since the first announcement of their result by Gröchenig and Lyubarskii:
Can the sufficiency condition be extended beyond the lattice setting? The answer is NO!
This paper provides explicit examples of Gabor systems with the first and third Hermite functions and oversampling rates of 3 and 5, respectively, which are not frames. These are derived from the location of known zeros of their Zak transforms (the existence of further zeros is open). Additionally, we show the existence of a Gabor system with the second Hermite function and oversampling rate 5, and one with the third Hermite function and oversampling rate 7, which are not frames. Our approach is not entirely new: The Zak transform and its zeros have often been used to construct lattices which, together with a (class of) Hermite window(s), do not yield a frame, see, e.g., [15, Prop. 3.5], [19], [21]. The sets we construct still exhibit periodicities and symmetries, but are not lattices. In [18], results for Gabor systems with semi-regular lattices $\Gamma\times{\mathbb{Z}}\subset{\mathbb{R}}^{2}$ have been derived. Our examples do not generally show this tensor structure, but we give non-frame examples for the third Hermite function with semi-regular lattices of density 2, 3, and 4.
A related problem concerns the frame set structure for Gabor systems with the $n$-th Hermite functions and lattices of density $11<D(\Lambda)\leq n+$ 1 [14]. So far, restrictions for odd Hermite functions and rectangular lattices are known [23], which have been extended to general lattices [6]: The Gabor system with an odd Hermite function is not a frame if the lattice density satisfies $D(\Lambda)=\frac{N+1}{N}$, $N$ a positive integer. For even Hermite windows, there were no further restrictions than the density condition $D(\Lambda)>1$ until Lemvig provided the first exceptions [21]. More examples of Gabor systems with Hermite functions and rectangular lattices of density $11<D(\Lambda)\leq n+$ 1, followed [19]. For $h_{1}$, it could still hold that the restriction $D(\Lambda)>\frac{N+1}{N}$, provided in [23], is the only exception [14]. The results in [8] broke a pattern occurring for Hermite functions of order up to 3, challenging a question posed by Lemvig [21]: Is the lattice density condition of Gröchenig and Lyubarskii sharp? This question remains open. If true, it may be answered by knowing the number and location of the zeros of the Zak transforms of generalized Hermite functions.
The author thanks the first referee for providing the following remarks and consequences: Our answer may be surprising. At least from a complex analysis viewpoint hints were making it reasonable to believe that a lower Beurling density $D^{-}(\Gamma)>n+1$ could be sufficient for separated points sets $\Gamma$ to form a Gabor frame with the $n$-th Hermite function (cf. [4] for the case of multiplicity sampling with Hermite functions). Denoting the upper Beurling density of $\Gamma$ by $D^{+}(\Gamma)$, it can be shown that a Gabor system with Hermite window is a Riesz basis if $D^{+}(\Gamma)<1/(n+1)$. Using the Weierstrass sigma function, this result was established for lattices in [1, Thm. 1]. However, as in [26], the result relies on the growth of the sigma function and can be generalized using a non-uniform version. In the case of a lattice $\Lambda\subset{\mathbb{R}}^{2}$, we have $D^{+}(\Lambda)=D^{-}(\Lambda)=D(\Lambda)$, which is called the density of the lattice. Duality results (Ron-Shen [24], Wexler-Raz [27]) yield equivalent conditions between the Riesz basis property and the frame property: A Gabor system is a frame if and only if the dual system with lattice $\Lambda^{\circ}=D(\Lambda)^{1/2}\Lambda$ is a Riesz basis. Thus, we not only show that the sufficient density result cannot be extended beyond the lattice setting, but our results also prohibit non-uniform versions of the Wexler-Raz or Ron-Shen duality principles. Otherwise, the density result could be deduced from the Riesz basis property.
Outline of the paper. We clarify the notation in § 2: In § 2.1 we introduce our normalization for Hermite functions, which is the same as in the book of Folland [11]. Time-frequency shifts and multi-window Gabor systems are introduced in § 2.2, following the book of Gröchenig [13]. Also, the notion of a periodic structure (the union of finitely many shifted copies of one lattice) is discussed there. We then present the (parameter-free) Zak transform in § 2.3, which maps functions from ${\mathbb{R}}$ to quasi-periodic functions with period lattice ${\mathbb{Z}}^{2}$. This is again covered by the book of Gröchenig [13]. The reader familiar with the concepts and notation may directly jump to the results in § 3. Finally, in § 4 we pose some open questions on the topic which should be understood as lines of research to advance the status of research. The author thanks the referees for the encouragement to include a discussion on open problems and Hermite windows of low order.
@@ 2. Notation
In most parts of this article, we will follow the notation from the book of Gröchenig [13]. The Hilbert space of square-integrable functions on the real line is denoted by $L^{2}({\mathbb{R}})$. For technical reasons, it will sometimes be necessary to restrict ourselves to a suitable dense subspace, in which case we will also write that a function is suitable. The statement is then valid for the Hilbert space by the density of the suitable function space, considering that some results will only hold almost everywhere and not pointwise.
We say that a function is suitable if it is in the modulation space $M^{1}({\mathbb{R}})$, also known as the Feichtinger algebra $S_{0}({\mathbb{R}})$ [9]. In some situations, it would suffice to pick a function from the (slightly larger) Wiener space of continuous functions $W_{0}({\mathbb{R}})$ with Fourier transform also contained in $W_{0}({\mathbb{R}})$, but to avoid confusion, we only consider $M^{1}({\mathbb{R}})$. As our results will be for Hermite functions, which are contained in the Schwartz space $\mathcal{S}({\mathbb{R}})$, the reader not familiar with $M^{1}({\mathbb{R}})$ may consider functions in $\mathcal{S}({\mathbb{R}})\subset M^{1}({\mathbb{R}})$ as suitable.
@@
Moreover, we will use the unitary dilation operator $\mathcal{D}_{a}$, $a>0$, which acts by the rule
@@ 2.1. Hermite functions
We consider Hermite functions as defined in [11]. The $n$-th order Hermite function is the $n$-th Hermite polynomial $H_{n}$ (subject to our normalization) multiplied by a Gaussian function $\phi(t)=e^{-\pi t^{2}}$. More precisely, we have
@@ A Gabor system in $L^{2}({\mathbb{R}})$ is a function system of the type
(2.8) $\pi(z)=M_{\omega}T_{x},\ z=(x,\omega)\quad T_{x}f(t)=f(t-x),\quad M_{\omega}f(t)=e^{2\pi i\omega t}f(t).$
@@
For more details, refer to [11, Chap. 1.3], [13, Chap. 9.1]. The important part for us here is that the composition of two time-frequency shifts is (up to a phase factor) again a time-frequency shift by the sum of the coordinates.
@@
We remark that we always assume $\Gamma$ to be separated (uniformly discrete), i.e.,
(2.13) $\inf_{\begin{subarray}{c}\gamma_{1},\gamma_{2}\in\Gamma\\ \gamma_{1}\neq\gamma_{2}\end{subarray}}(|\gamma_{1}-\gamma_{2}|)\geq\delta>0.$
In this case, for $g\in M^{1}({\mathbb{R}})$, the upper frame bound is always finite, and we only need to care about the strict positivity of the lower frame bound. In the sequel, we will consider periodic configurations. The points then, indeed, have a minimal distance from each other.
@@
(2.14) $\mathfrak{G}(\{g_{m}\}_{m=1}^{M},\Gamma)=\bigcup_{m=1}^{M}\mathcal{G}(g_{m},\Gamma)$
@@
(2.15) $S_{\mathfrak{G}(\{g_{m}\}_{m=1}^{M},\Gamma)}=\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},\Gamma)}=\sum_{m=1}^{M}\sum_{\gamma\in\Gamma}\langle f,\pi(\gamma)g_{m}\rangle\,\pi(\gamma)g_{m}.$
The windows $g_{m}$ may be completely different, but we are interested in the case where they are shifted copies $\pi(z_{m})g$ of a single function $g$. A direct computation shows that the multi-window frame operator is also the associated frame operator of a Gabor system over a periodic structure $\Gamma(\Lambda,\{z_{m}\}_{m=1}^{M})$, $\Lambda$ a lattice, with window $g$. The periodic structure is
(2.16) $\Gamma(\Lambda,\{z_{m}\}_{m=1}^{M})=\bigcup\limits_{m=1}^{M}(\Lambda+z_{m})$
To simplify notation, we set $\Gamma_{M}(\Lambda)=\Gamma(\Lambda,\{z_{m}\}_{m=1}^{M})$, keeping in mind that we still need to choose the set $\{z_{1},\ldots,z_{M}\}$. While $\Gamma_{M}(\Lambda)$ is still $\Lambda$-periodic, it is not necessarily a lattice. The density, or (over-)sampling rate, of a periodic structure is given by
(2.17) $D(\Gamma_{M}(\Lambda))=M\cdot D(\Lambda)=\frac{M}{{\rm vol}({\mathbb{R}}^{2}/\Lambda)}.$
Assume $g$ is suitable (so the following series converges unconditionally), then
(2.18) $S_{\mathfrak{G}(\{\pi(z_{m})g\}_{m=1}^{M},\Lambda)}f$ $=\sum_{m=1}^{M}\sum_{\lambda\in\Lambda}\langle f,\pi(\lambda)\pi(z_{m})g\rangle\,\pi(\lambda)\pi(z_{m})g$
@@
The phase factor, which appears due to (2.9), also appears as a complex conjugate in the above calculation and cancels. The last detail we collect is the unitary equivalence of the Gabor systems $\mathcal{G}(g,\Gamma)$ and $\mathcal{G}(\mathcal{D}_{a}g,D_{a}\Gamma)$ (see [7], [12]), where
@@
First, for $z\in{\mathbb{R}}^{2}$, we note that $\mathcal{D}_{a}\,\pi(z)\,\mathcal{D}_{a}^{-1}=\pi(D_{a}z)$, which can be checked by a simple calculation with a suitable function. From this, we get
@@
In the above computations, we used the unconditional convergence of the series (because $g$ is suitable, so we have a Bessel system, i.e., a finite upper frame bound) and the fact that $\mathcal{D}_{a}$ is unitary. In particular, the two Gabor systems have the same optimal frame bounds.
@@ 2.3. The Zak transform
Thus, $\mathcal{Z}f(x,\omega)$ is quasi-periodic with respect to the integer lattice ${\mathbb{Z}}^{2}$ and completely determined by the values in $[0,1)^{2}$. Next, we note that a time-frequency shift of the function basically results in a shift of the Zak transform. Due to (2.9), we obtain
@@
the phase factor in (2.27) appears then due to the commutation relation (2.9). We also note that the Zak transform extends to a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}([0,1)^{2})$. We refer to [13, Thm. 8.2.3] for the details of this fact.
We close the survey on the Zak transform by listing what we call the trivial zeros. They occur due to the parity of $f$. For the details see [20, § 5] or [19, § 4]. We have (see Figure 1):
  • For a suitable $f$, if $f(t)=f(-t)$, so $f$ is an even function, then
@@
Here, the condition that $f$ is suitable can be weakened to $f\in W_{0}({\mathbb{R}})$ but cannot be extended to $f\in L^{2}({\mathbb{R}})$. It implies that $\mathcal{Z}f$ is continuous (see [13, Prop. 8.2.1 (c)]). A remarkable property of the Zak transform is that, if $\mathcal{Z}f$ is continuous, then it has (at least) a zero in $[0,1)^{2}$. The indicator function $\chi_{[0,1]}$ is not suitable and not in $W_{0}({\mathbb{R}})$. It is not continuous, hence, not in $M^{1}({\mathbb{R}})\subset W_{0}({\mathbb{R}})\cap\mathcal{F}W_{0}({\mathbb{R}})$, and $|\mathcal{Z}\chi_{[0,1]}|=1$ almost everywhere.
@@ 3. Results
We now present a result of how the zeros of $\mathcal{Z}g$ determine the frame property of integer oversampled Gabor systems. The new aspect is that we do not insist that the underlying set is a lattice. Thus, the result belongs to the study of non-uniform Gabor systems. We refer to [2], [5], [10], [17] to name just a few results in this direction.
@@
For $g\in M^{1}({\mathbb{R}})$, let $\{z_{1},\ldots,z_{M}\}\subset[0,1)^{2}$ be distinct zeros of its Zak transform, i.e., $\mathcal{Z}g(z_{m})=0$, $m=1,\ldots,M$. Then, the multi-window Gabor system $\mathfrak{G}(\{\pi(z_{m})g\}_{m=1}^{M},{\mathbb{Z}}^{2})$ is not a frame for $L^{2}({\mathbb{R}})$. Equivalently, we have that the Gabor system $\mathcal{G}\left(g,\bigcup_{m=1}^{M}({\mathbb{Z}}^{2}+z_{m})\right)$ is not a frame for $L^{2}({\mathbb{R}})$.
@@
To keep this article self-contained, we repeat a standard proof, as given in the book of Gröchenig [13, Chap. 8.3], that the Zak transform diagonalizes the frame operator. More precisely, for a multi-window Gabor system we will show (cf. [13, eq. (8.18)]) that
@@
(3.2) $\mathcal{Z}\left(\sum_{m=1}^{M}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right)=\sum_{m=1}^{M}\left(\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f\right).$
@@
(3.5) $=\left(\sum_{(k,l)\in{\mathbb{Z}}^{2}}\langle\mathcal{Z}f,\mathcal{Z}(M_{l}T_{k}g_{m})\rangle_{L^{2}([0,1]^{2})}\,e^{2\pi i(k\omega-lx)}\right)\,\mathcal{Z}g_{m}(x,\omega).$
We have a look at the inner product. Using (3.3), we obtain
@@
These are Fourier coefficients of $\mathcal{Z}f\overline{\mathcal{Z}g_{m}}$, and the whole expression in the parentheses in (3.5) is simply a Fourier series of this expression. Thus,
(3.7) $\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}f=\left(\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}\right)\mathcal{Z}f\quad\Longleftrightarrow\quad\sum_{m=1}^{M}\mathcal{Z}S_{\mathcal{G}(g_{m},{\mathbb{Z}}^{2})}\mathcal{Z}^{-1}=\sum_{m=1}^{M}|\mathcal{Z}g_{m}|^{2}.$
@@
(3.9) $\sum_{m=1}^{M}|\mathcal{Z}(\pi(z_{m})g)(z)|^{2}=\sum_{m=1}^{M}|\mathcal{Z}(g)(z+z_{m})|^{2},\quad z=(x,\omega).$
@@
forcing the lower frame bound to vanish, and the result is proved. ∎
The methods in the proof of Theorem 3.1 may be considered standard in time-frequency analysis. The novel aspect is that we apply it to a periodic structure, which can be decomposed into finitely many shifted copies of a lattice. So far, the method has only been applied to lattices that could be separated into shifts of a coarser lattice. In the same spirit, our approach can be extended to periodic structures of rational density by using techniques as presented by Zibulski and Zevii [29] (see also [13, Chap. 8.3] or [23]).
Note that the result also holds for $\mathcal{G}(g,\bigcup_{m=1}^{M}({\mathbb{Z}}^{2}+z_{m}+\xi))$, i.e., if we consider a global shift by $\xi\in{\mathbb{R}}^{2}$, as the location of all zeros in (3.10) is shifted simultaneously. As $\pi$ is unitary, the invariance under global shifts generally reflects itself in the frame inequality (2.11):
(3.11) $\sum_{\gamma\in\Gamma+\xi}|\langle f,\pi(\gamma)g\rangle|^{2}=\sum_{\gamma\in\Gamma}|\langle f,\pi(\gamma+\xi)g\rangle|^{2}=\sum_{\gamma\in\Gamma}|\langle\pi^{-1}(\xi)f,\pi(\gamma)g\rangle|^{2},\quad\forall f\in L^{2}({\mathbb{R}}).$
With these arguments, all results for Gabor systems hold true for global shifts in the sequel.
At this point, we remark that several versions of the Zak transform exist. For example, in [19] or [21] the following version is used (re-call $\mathcal{D}_{a}$ from (2.3))
@@
This can be used to study Gabor systems of the form $\mathcal{G}(\mathcal{D}_{a}^{-1}g,{\mathbb{Z}}^{2})$. By unitary equivalence, this is the same as studying properties of $\mathcal{G}(g,D_{a}{\mathbb{Z}}^{2})$, so a Gabor system with window g over the rectangular lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. On the other hand, there is the following version, which is, for example, presented in [20]. Setting $z=(x,\omega)\in{\mathbb{R}}^{2}$, we can write this version as
(3.13) $\mathcal{Z}_{a}f(z)$ $=\sqrt{a}\sum_{k\in{\mathbb{Z}}}f(ak-x)e^{2\pi iak\omega}=\sum_{k\in{\mathbb{Z}}}\mathcal{D}_{a}^{-1}\pi(z)f(k)=\sum_{k\in{\mathbb{Z}}}\pi(D_{a}^{-1}z)\mathcal{D}_{a}^{-1}f(k)$
@@
The essential properties from $\mathcal{Z}$ carry over to $\mathcal{Z}_{a}$, but the main difference is that $\mathcal{Z}_{a}$ is quasi-periodic with respect to the lattice $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$. Hence, we have that $\mathcal{Z}_{a}$ is determined by its values on $[0,a)\times[0,\frac{1}{a})$ and it is a unitary operator from $L^{2}({\mathbb{R}})$ onto $L^{2}(D_{a}[0,1]^{2})$. This version of the Zak transform, combined with the techniques presented in the proof above, has long been used to study Gabor systems over rectangular lattices of the form $a{\mathbb{Z}}\times b{\mathbb{Z}}$ with $(ab)^{-1}\in{\mathbb{N}}$. All results in this article carry over to $\mathcal{Z}_{a}$ and $a{\mathbb{Z}}\times\frac{1}{a}{\mathbb{Z}}$ with minor adjustments. This technique has also been expanded to the case that $(ab)^{-1}=\frac{p}{q}\in{\mathbb{Q}}$, leading to determining whether certain $q\times q$ matrices have full rank (see [13, Chap. 8.3]). A rectangular lattice as an index set is quite restrictive, as the shifts $\pi(\xi_{n})$ must possess very restricted symmetries. This is discussed to some extent in [19, § 3].
The freedom of allowing arbitrary periodic configurations leads to the following result.
Proposition 3.2.
@@
The fact that $g\in M^{1}({\mathbb{R}})$ guarantees that we have a finite upper frame bound. Thus, similar to (3.10), in order to have a frame, we need to ensure that
(3.16) $\min_{z\in[0,1)^{2}}\sum_{m=1}^{M}|\mathcal{Z}g(z+z_{m})|^{2}>0.$
As the zero set of $\mathcal{Z}g$ is discrete and finite by assumption, the probability that $K$ i.i.d. randomly picked points from $[0,1)^{2}$ all belong to (a shift of) the zero set of $\mathcal{Z}g$ is zero. ∎
A consequence of Theorem 3.1 and the results of Gröchenig and Lysubarskii [15][16] concerns the structure of the location of the zeros of Zak transforms of Hermite functions.
@@
[…]